Prove that if f and g are Riemann integrable on [a, b], then f · g and f + g are Riemann integrable on [a, b].


1
Expert's answer
2022-03-15T19:31:28-0400



Let Df, Dbe set of all discontinuations of f and g.

f+g is integrable iff,

"Df\bigcap Dg"

has a measure zero

We have,

"Df\bigcap Dg \subset" "Df" giving the set on left having a measure zero

Hence f+g is integrable


If f,g are both Riemann integrable then f.g is also integrable

this is proved by proving that fis integrable

"f(g)=\frac{1}{2}((f+g)^2-f^2-g^2)"


"U(f^2,P)-L(f^2,P)
=\sum(Mi^2-m
i^2)\Delta xi 

<2T\frac{\epsilon}{2T}=\epsilon"

Here, "Sup f(x) of f^2" is integrable and therefore f.g is integrable.